OpenStudy (anonymous):

when finding the dirivitive of f(x,y) =2x^2y^3 using the fuction at (a,b) and holding y constant.. you get g(x)=f(x,b) =2x^2b^3 .... THEN rate of change of the function at (a,b) if we hold y fixed and allow x to vary... g'(a)=4ab^4 ....... where does this part come from, how do you get to 4ab^4 from 2x^2b^3?

zepdrix (zepdrix):

Hmm shouldn't you get 4ab^3? Is that a typo by chance? Maybe I'm mistaken :) Lemme look at the problem again.. hmm

zepdrix (zepdrix):

Hmm I guess this one has me stumped. :\ @Hero @hartnn @experimentX

hartnn (hartnn):

Holding 'y' constant means partially differentiating w.r.t 'x' \(\large \frac{\partial}{\partial x}(2x^2y^3)=4xy^3 \\then, g(x)=f'(x,b)=4xb^3\) in the question it appears that derivative is not taken at all, even if its mentioned..

hartnn (hartnn):

i would request the asker to post the screenshot of actual question.

OpenStudy (anonymous):

this is what i am looking at http://tutorial.math.lamar.edu/Classes/CalcIII/PartialDerivatives.aspx where does the 4ab(cubed) come from?

zepdrix (zepdrix):

So it was a typo? Darn I wish you woulda mentioned that :) lol

hartnn (hartnn):

right, so since 'y' was constant, it was taken as b, (y=b) hence, the function f is now only function of x = g(x) = \(2x^2b^3\) Now, to get rate of change, we differentiate w.r.t x \(g'(x)=4xb^3\) to find rate of cahnge AT x=a we just plug in x=a in that. \(g'(a)=4ab^3\) got this ?

hartnn (hartnn):

another way to represent g'(a) \(g'(a)=4ab^3=f_x(a,b)\)

OpenStudy (anonymous):

I am having trouble doing anything on this site, it is lagging like hell.. have a few questions regarding partial derivitives.. if you guys can shoot me an email that would be awesome.. fivestarsmi@gmail.com

OpenStudy (anonymous):

everything you post comes up as math processing error, I cant even view it =/

hartnn (hartnn):

let me post you an image...of my comment... can u view this ? and understand ?

OpenStudy (anonymous):

I understand everything, except where the 4xb^3 comes from

OpenStudy (anonymous):

how do you get from 2x^2b^3 to 4xb^3?

hartnn (hartnn):

the derivative part ?? derivative of x^n is nx^(n-1) so, derivative of 2x^2 will be 2*(2x^{2-1}) = 4x and since b^3 is constant, it can be taken out of derivative. so, (2x^2b^3)' = 4xb^3

hartnn (hartnn):

(2x^2b^3)' =2b^3(x^2)'=2b^3(2x)= 4xb^3

hartnn (hartnn):

any further doubts ?

OpenStudy (anonymous):

is there a name for this formula to find the dirivitive? the derivative of x^n is nx^(n-1)

OpenStudy (anonymous):

is there a name for this formula to find the dirivitive? the derivative of x^n is nx^(n-1)

hartnn (hartnn):

no... i don't think there is any name to it...it helps in differentiating any polynomial....

OpenStudy (anonymous):

how do you describe differntiating in laymans terms lol

hartnn (hartnn):

if a function f(x) is increased by very small amount say, 'h' (small means h->0) then differentiation of 'x' or rate of change of the function f(x) = f'(x) will be just the ratio of change in function / small change h. and the mathematical formula will be lim h->0 {f(x+h)-f(x)}/{h} i don't know whether this is layman's terms, but ask if you have any doubts in this....

OpenStudy (anonymous):

ok, in your first reply you have the partial differential symbol of the partcial dif of x... does that just mean to apply the x^n is nx^(n-1) formula? ... how is that read?

hartnn (hartnn):

the question had x^2 in it, thats why i used x^n formula. i read it as 'daaba by daaba x' but different people read it differently. technically its partial derivative of function f w.r.t x

OpenStudy (anonymous):

qhat dould it mean if it had daaba z by dadda x

OpenStudy (anonymous):

for instance in the navier-stokes formula.. it has daaba v by dabba t ... is the v for viscosity or vector?

OpenStudy (anonymous):

v is for flow velocity..

hartnn (hartnn):

z can be function of many variable, like x,y,p,q....and so on now partial derivative of z w.r.t x means treat all the variables other than 'x' as CONSTANT. now z will only be function of 'x' and just diff. w.r.t x normally.

OpenStudy (anonymous):

so daaba v by daaba t .. the partial dif of the flow velocity over the partial dif of t? (time?)

hartnn (hartnn):

the partial dif of the flow velocity with respect to 't' means only 't' is variable, all other letters are constant

OpenStudy (anonymous):

Cartesian coordinates Writing the vector equation explicitly.... what does this mean, when you have several partial dirivities in an equation/.

hartnn (hartnn):

ok, 'u' is the function of t,x,y,z daaba u by daaba t implies derivative of 'u' treating x,y,z as constants and 't' as variable. daaba u by daaba x implies derivative of 'u' treating t,y,z as constants and 'x' as variable. daaba u by daaba y implies derivative of 'u' treating x,t,z as constants and 'y' as variable. daaba u by daaba z implies derivative of 'u' treating x,y,t as constants and 'z' as variable.

OpenStudy (anonymous):

what does smooth mean?

hartnn (hartnn):

smooth means not rapidly changing.....

OpenStudy (anonymous):

s = (x, y, z) \[\in\] ℝ3 does \[\in\]just means in the 3d space?

hartnn (hartnn):

∈ means 'belongs to' R3 means 3 dimentional space.